OpenStudy (anonymous):

The decay of uranium is modeled by D=D1 x 2^-kt. If it takes 6 years for the mass of uranium to halve, find the percentage remaining after 2 years

OpenStudy (anonymous):

"k" in the equation is called the decay constant of the sample i.e., reciprocal the time taken for half the initial amount (no matter how much iniial, but half of it) to decay. k = 1 / (half-life time) the question asks you to find % remaining = (initial amount - lost) *100 / (initial amount) = (D1-D)*100/D1 =(1 - D/D1)*100 given t=2

OpenStudy (anonymous):

I didn't get what you were trying to explain

OpenStudy (anonymous):

decay const : k = 1 / (time to decay to half) = ?

OpenStudy (anonymous):

I'll write it down on paper to see if it makes sense

OpenStudy (anonymous):

\[k=\frac{1}{t_{1/2}}\] \(t_{1/2}\) is given in the problem

OpenStudy (anonymous):

I still can't get it

OpenStudy (anonymous):

I mean the whole formula you mentioned above

OpenStudy (anonymous):

you are given\[D=D_1e^{-kt}\] correct?

OpenStudy (anonymous):

How to find the percentage and how can I find D and D1

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

It's 2 instead of e

OpenStudy (anonymous):

D1 = total amount initially present D = amount left after time t

OpenStudy (anonymous):

correct, "2" instead of "e"

OpenStudy (anonymous):

agree? the question is asking you what % is left after time t

OpenStudy (anonymous):

Now I'm beginning to understand

OpenStudy (anonymous):

so, they are asking for \({D\over D_1}\times100\)

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

so,\[\%left = 2^{-kt},\qquad k={1\over t_{1/2}}\] \(t_{1/2}\) is the time taken to reduce to half the amount

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

And we have the value of t which is 2

OpenStudy (anonymous):

you may proceed from here. 1. find k using given \(t_{1/2}\) (using the equation on the right) 2. use that k and given "t" (using equation to the left) that is your answer

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Would k =1 ?

OpenStudy (anonymous):

no, step "1"

OpenStudy (anonymous):

Sorry 1/2?

OpenStudy (anonymous):

what is \(t_{1/2}\)?

OpenStudy (anonymous):

Isn't it 2?

OpenStudy (anonymous):

how? it is provided in the problem

OpenStudy (anonymous):

"6"

OpenStudy (anonymous):

what happens in "6" years?

OpenStudy (anonymous):

The uranium halves

OpenStudy (anonymous):

that is \(t_{1/2}\) =half life time

OpenStudy (anonymous):

So k is 1/6?

OpenStudy (anonymous):

half life = time taken for the amount of material to reduce to half-its initial amount

OpenStudy (anonymous):

yes k=1/6

OpenStudy (anonymous):

:) sorry man I'm giving you a hard time

OpenStudy (anonymous):

thats kool.

OpenStudy (anonymous):

Now to find the % I have to insert k (1/6) and t (2) into the equation?

OpenStudy (anonymous):

that's it!

OpenStudy (anonymous):

Then I multiply by 100

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Could you please show me how just to make sure

OpenStudy (anonymous):

\[\%=2^{-{1\over6}\times2}\times100=2^{-1\over3}\times100=?\]

OpenStudy (anonymous):

79%

OpenStudy (anonymous):

I've got one more thing to ask you about

OpenStudy (anonymous):

yes.. sounds reasonable.

OpenStudy (anonymous):

You wrote previously above that % = (D1-D) x 100/D1

OpenStudy (anonymous):

Where did you deduct it from?

OpenStudy (anonymous):

yes.. my bad.. I thought D represented amount "lost" but D represents amount "left"

OpenStudy (anonymous):

Is that equation still correct ?

OpenStudy (anonymous):

the question does not tell us about this abiguity

OpenStudy (anonymous):

Is that another way of finding % if we were given D and D1?

OpenStudy (anonymous):

those are the two ways.

OpenStudy (anonymous):

based on the interpretation of the words of the question

OpenStudy (anonymous):

Is there any other way other than these two ways?

OpenStudy (anonymous):

Anyways thanks a lot for your help. I really appreciate it.

OpenStudy (anonymous):

you are welcome